OpenStudy (anonymous):

I need help Casey was building a rectangular pen for his pigs. He has 62 feet of fencing. The length of his pen is 9 feet longer than the width. Write and solve an equation to find the dimensions of the pen.

OpenStudy (campbell_st):

let the width by w so the length is w + 9 so the perimeter is 62 = 2w + 2l or 62 = 2w + 2(w + 9) simplify the equation and then solve for w

OpenStudy (anonymous):

11 feet by 20 feet.

OpenStudy (anonymous):

2x+2(x+9)=62 OR 4x+18=62 The dim. of the pen are 11 feet by 20 feet